OpenStudy (anonymous):

Find the area of each circle. Round to the nearest tenth 1. a circle with a radius of 6 yards . . 2. a circle with a diameter of 18 millimeters

OpenStudy (hoblos):

area = (pi)r^2

OpenStudy (hoblos):

so for the first area = (pi)6^2 = 113.1 yd^2 for the second r=18/2 =9 area=(pi)r^2 = 254.5 mm^2

OpenStudy (anonymous):

Helllppppp. Find the area of each regular polygon. Round to the nearest tenths. 1. A pentagon with a perimeter of 45 feet . 2. a hexagon with a side length of 4 inches. 3. a nonagon with a side length of 8 meters 4. a triangle with perimeter of 54 centimeters